• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name Yi Qiang

    Log entry time 08:02:54 on June21,2004

    Entry number 126482

    keyword=owl shift summary

    00:00 Start to ramp up Right Septum and left Septum is still
    ramping. Once left arm is ready, we'll ask for 30uA beam
    and take data from left arm with right septum ramping.
    00:30 Right septum rampped much faster than left, we can expect
    to start run with both arms ready.
    00:40 Both Arms are ready!!!! Call MCC for beam @ 30uA
    01:00 Change HallC IA setpoint to 5.505
    01:17 Adjust HallC IA setpoint to 5.650, now HallC BCM asym
    value is arount -80
    02:00 Lost control of Septa, try to reboot PLC
    02:20 Reboot again w/o sucess, page Todd (tech on call) and
    call MCC for control access
    03:00 Todd Ewing came
    03:30 Todd went down hall to reset the computers
    04:50 Both Septa are in Alive Status
    05:00 Left arm Q1 tripped, called Todd, they will fix it
    05:15 Q1 is back, it looks related with septa rebooting
    06:50 Septa are back!!!!!! Thanks Mark and Todd!!! Begin rampping
    07:28 Right Septum Rampped up, start a prodcution run

    ===============================================================
    Run Summary
    2321 00:46-01:39 production 20cm He
    2322 01:40-stopped for septa problem junk 20cm He
    2323 02:24-02:25 septa test junk
    2324 06:54-07:28 test junk
    2325 07:29- production 20cm He Right Only

    ===============================================================
    Spec Run
    1153 01:39-01:41 Raster check